OpenStudy (anonymous):

A piece of chalk is tossed vertically upward by Prof. Schwarzenegger and hits the ceiling 98 feet above with a BANG. (Neglect the effects of air resistance. Round your answers to two decimal places.) (a) What is the minimum speed the piece of chalk must have been traveling to enable it to hit the ceiling? ft/s (b) Assuming that Prof. Schwarzenegger in fact tossed the piece of chalk up at 98 ft/s, how fast was it moving when it struck the ceiling? ft/s (c) Assuming that Prof. Schwarzenegger tossed the chalk up at 98 ft/s, and that it recoils from the ceiling with the same speed it had at the A piece of chalk is tossed vertically upward by Prof. Schwarzenegger and hits the ceiling 98 feet above with a BANG. (Neglect the effects of air resistance. Round your answers to two decimal places.) (a) What is the minimum speed the piece of chalk must have been traveling to enable it to hit the ceiling? ft/s (b) Assuming that Prof. Schwarzenegger in fact tossed the piece of chalk up at 98 ft/s, how fast was it moving when it struck the ceiling? ft/s (c) Assuming that Prof. Schwarzenegger tossed the chalk up at 98 ft/s, and that it recoils from the ceiling with the same speed it had at the

OpenStudy (amistre64):

-16t^2 + Vt = 98 -16t^2 + Vt -98 = 0

OpenStudy (amistre64):

hmmmm

OpenStudy (anonymous):

is that one of the answers?

OpenStudy (amistre64):

this is a study group right?

OpenStudy (amistre64):

its a thought i have as to what an answer might be, but i havent got to the end of the thought yet

OpenStudy (amistre64):

the derivative of that position formula tells us the speed, the velocity, at any given moment

OpenStudy (amistre64):

so when it is at 0 we know that is the least amount of velocity needed to hit the ceiling

OpenStudy (anonymous):

im sorry i just have no idea what to do for this problem any ideas help

OpenStudy (amistre64):

-16t^2 + Vt -98 = 0 -32t + V = 0 when t = V/32 .... so we could sub that back in to "t" to see if that helps out with the V

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=simplify+-16%28V%2F32%29^2+%2B+V%28V%2F32%29+-98 the wolf simplifies nicely :)

OpenStudy (anonymous):

so which answer goes with a, b, and c? hahaha

OpenStudy (amistre64):

when we solve this for V we only got one good option: 56sqrt(2) i beleive

OpenStudy (amistre64):

I believe a is 56sqrt(2); but id have to dbl chk it with the wolf

OpenStudy (anonymous):

the 56th square root of 2?

OpenStudy (amistre64):

no, 56 times the sqrt(2)

OpenStudy (amistre64):

b and c are physics parts that i am not familiar with

OpenStudy (amistre64):

good luck with the rest of it, or if you have any questions about part a let me know

OpenStudy (anonymous):

wow a was right thank you so much

OpenStudy (amistre64):

yay!! lol

OpenStudy (anonymous):

haha thanks so much

OpenStudy (anonymous):

can you solve it for 105 feet by any chance?